a prism with critical angle 30 immersed in medium of critical angle 45, what is the final critical angle of prism?
 

Grade:12th pass

1 Answers

Khimraj
7 years ago
As sinC = 1/n so final critical angle is SinC = ref index of medium/ref index of prism = sin critical angle of prism/ sin of critical angle of medium = (1/2)/(1√2) = 1/√2 . So new critical angle is 45°.Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer.